can you please explain this to me ok r/3=x/x
Answer by stanbon(75887)

can you please explain this to me ok r/3=x/x
-----
If the "x" in the right side is not zero, the
right side is equal to "1".
---
Solve: r/3 = 1
---
Multiply both sides by 3 to solve for "r":
r = 3
============
Cheers,
Stan H.